# Sparsity-Promoting Reachability Analysis and Optimization of Constrained Zonotopes

### v2

@@ III-B Numerical Example: Reachable Sets of a Second-Order System
Fig. 2 shows the reachable sets $\mathcal{X}_{k}$ for this system for $k\in\{0,...,N\}$ where $N=15$. The sparsity patterns for the resulting constrained zonotope generator and constraint matrices $G$ and $A$ used to represent $\mathcal{X}_{N}$ are shown in Fig. 3. The number of non-zero elements in $G$ and $A$ are given explicitly in Table I. Reachability calculations using Eq. (17) result in the constrained zonotope with the fewest non-zero elements in both $G$ and $A$. Further, this approach is tied with (10) for producing a constrained zonotope with the fewest number of generators $n_{G}$ and constraints $n_{C}$.
@@
Figure 3: Sparsity plots for matrices $G$ and $A$ where $\mathcal{X}_{N}=\left\langle G,\mathbf{c},A,\mathbf{b}\right\rangle$.
TABLE I: Number of non-zero elements in the generator matrix $G$ and constraint matrix AA for $\mathcal{X}_{N}$.
@@
$\operatorname*{\mathrm{nnz}}{(G)}$ 33 5 2
$\operatorname*{\mathrm{nnz}}{(A)}$ 315 237 105
@@ IV-A ADMM Formulation
it is clear that (21) is a QP with constraints $\bm{\xi}\in\mathcal{A}\cap\mathcal{B}$ where $\mathcal{A}=\left\{\bm{\xi}\in^{n_{G}}\middle|A\bm{\xi}=\mathbf{b}\right\}$ is an affine set and $\mathcal{B}=\left\{\bm{\xi}\in^{n_{G}}\middle|-\mathbf{1}\leq\bm{\xi}\leq\mathbf{1}\right\}$ is a box. Using interval notation, we will alternately write $\mathcal{B}$ as $\bm{\beta}]$.
@@
Define the matrix $M$ to be
$M=\begin{bmatrix}\tilde{P}+\rho I&A^{T}\\ A&0\end{bmatrix}\;$. (24)
The ADMM iterations (5) then reduce to
@@
where $\min(\cdot,\cdot)$ and $\max(\cdot,\cdot)$ denote element-wise operations.
@@
If Assumption 1 does not hold, then redundant constraints can be removed as in [18, Alg. 1].
@@
The ADMM formulation in (25) exploits the structure of the constrained zonotope set representation. As an illustrative example, consider the case where $\mathcal{Z}$ is a centrally symmetric hexagon $\mathcal{Z}=\left\langle G=\begin{bmatrix}\mathbf{g}_{1}&\mathbf{g}_{2}&\mathbf{g}_{3}\end{bmatrix},\mathbf{c},[],[]\right\rangle$ and $P=I$. The matrix $M$ then becomes $M=G^{T}G+\rho I$ which has up to $9$ non-zero elements. Representing the same set using H-rep polytopes requires $6$ inequality constraints $\mathbf{a}_{i}^{T}\mathbf{x}\leq b_{i}$. The addition of $6$ slack variables $s_{i}>0$ transforms the inequality constraints into equality constraints such that the problem can be solved via a more general version of the ADMM formulation given in (25) (see [37, 38, 39]). The resulting $A$ matrix has up to $18$ non-zero elements in this case and $P+\rho I$ has $8$ non-zero elements due to the addition of the slack variables, so $M$ will have up to $44$ non-zero elements using H-rep.
Note that $M$ will not have fewer non-zero elements using a constrained zonotope representation $\mathcal{Z}$ vs. H-rep in general. In practice, $M$ has fewer non-zero elements when (1) sparsity-promoting operations are used to compute $\mathcal{Z}$ (Sec. III), and (2) the number of equality constraints in $\mathcal{Z}$ is minimized (e.g., by using operations on zonotopes to build $\mathcal{Z}$ when possible).
The ADMM formulation in (25) is more concise for optimizing over constrained zonotopes than other ADMM formulations. For example, the widely used ADMM solver OSQP [24] solves QPs with constraints of the form $C\mathbf{x}=\mathbf{z}$, $\mathbf{z}_{l}\leq\mathbf{z}\leq\mathbf{z}_{u}$. To optimize over $\mathcal{Z}$ using this form requires $C=\begin{bmatrix}A^{T}&I^{T}\end{bmatrix}^{T}$, which adds $2n_{G}$ non-zero elements to the linear system that OSQP must solve and adds $n_{C}$ projection operations.
Combining constrained zonotopes with high sparsity and the ADMM iterations (25) results in an approach to solving QPs that scales well to large problem sizes, as will be shown in Sec. VI-B.
@@ IV-B Infeasibility Detection
In the following theorem, we propose an infeasibility certificate for (23) that is exact within machine precision and does not require any user-defined tolerances.
@@
Problem (23) is infeasible if $\exists\mathbf{v}\in\text{span}(A^{T})$ such that $\mathbf{v}^{T}\tilde{\bm{\xi}}\notin\mathbf{v}^{T}[\bm{\beta}]$ where $\tilde{\bm{\xi}}$ is any point in $\mathcal{A}$.
@@
If (23) is infeasible and $Z^{T}\tilde{P}Z$ is positive definite where $Z$ is a basis for the null space of AA, then the difference of the iterates $\bm{\zeta}_{k}-\bm{\xi}_{k}\rightarrow\mathbf{v}$ for $k\geq k^{\prime}$ with $k^{\prime}$ sufficiently large, where $\mathbf{v}$ is an infeasibility certificate for problem (23).
@@
To implement Thm. 1 for infeasibility detection, the difference of the iterates $\bm{\zeta}_{k}-\bm{\xi}_{k}$ is projected onto the span of $A^{T}$ to generate a candidate infeasibility certificate vector $\mathbf{v}$. If $\mathbf{v}$ satisfies Thm. 1, then (23) is infeasible regardless of whether $k\geq k^{\prime}$ as defined in Cor. 1. In practice, this results in infeasibility certificates being generated in as few as $k=1$ iterations for many problems, as will be shown in Sec. VI-D. Empirically, infeasibility certificates are produced using this implementation for $Z^{T}\tilde{P}Z$ positive semi-definite as well.
@@ IV-C Interpretation of ADMM Iterations
Consider an MPC optimization problem for the system (9) for which the feasible set is given as
$\mathcal{Z}=\mathcal{X}_{0}\times\mathcal{U}\times\mathcal{X}_{1}\times\dots\times\mathcal{U}\times\mathcal{X}_{N}\;.$ (26)
@@
In [41], an MPC formulation using ADMM with early termination was presented that decoupled enforcement of system dynamics constraints from state and input feasibility constraints between ADMM steps. In that formulation, the equality constraints enforced in (5a) correspond to the system dynamics while feasibility constraints are enforced in (5b). This is analogous to building the MPC problem using (17) with zonotopes in our framework. As described above, using (10) with zonotopes to build the MPC problem would instead simultaneously enforce dynamics and input constraints with state constraints decoupled.
@@ V Applications
The following application examples demonstrate the utility of combining the sparse reachability operations presented in Sec. III with the ADMM algorithm presented in Sec. IV to solve control, estimation, and safety verification problems.
@@ V-A Model Predictive Control
$\mathbf{x}_{k},\mathbf{x}_{N}\in\mathcal{S}_{k}\;,\;\mathbf{u}_{k}\in\mathcal{U}\;,$ (27c)
@@
13:return  $\left(\mathcal{Z},P,\mathbf{q}\right)$
@@ V-B Moving Horizon Estimation and Set-Valued State Estimation
Assuming $\mathcal{S}$, $\mathcal{W}$, $\mathcal{V}$, and $\mathcal{X}_{-N}$ are zonotopes or constrained zonotopes, constrained zonotope reachability analysis can be used to solve the SVSE problem for (28) via the iterations
@@
The process and measurement noise covariance matrices are $Q$ and $R$, respectively. The covariance of the state estimate at time step k=Nk=-N is $P_{-N}$, and the prior state estimate for k=Nk=-N is $\hat{\mathbf{x}}_{-N}$. The set of possible states at time step k=Nk=-N is $\mathcal{X}_{-N}$, which is updated recursively between MHE iterations by applying (30).
Algorithm 3 shows how the MHE optimization problem (31) can be written in the form of (20) using constrained zonotopes and the set operations in (3). Line 5 implements (30) without projecting onto $\mathcal{X}_{k+1}$. The constrained zonotope $\mathcal{Z}$ representing the set of feasible solutions to (31) has the structure
$\mathcal{Z}=\mathcal{X}_{-N}\times\mathcal{W}\times\mathcal{X}_{-N+1}\times\dots\times\mathcal{W}\times\mathcal{X}_{0}\;,$ (32)
and as such the solution $\mathcal{X}_{0}$ to the SVSE problem at time step $k=0$ can be readily extracted via a linear map operation. After solving (20), a best state estimate $\hat{\mathbf{x}}_{0}\in\mathcal{X}_{0}$ can also be extracted.
@@ VI-A Implementation
All examples are implemented using the open-source toolbox ZonoOpt [^1] https://github.com/psu-PAC-Lab/ZonoOpt, which was developed as a contribution of this paper. This toolbox provides classes for zonotopes, constrained zonotopes, and hybrid zonotopes, implements the set operations in (3) (among others), and includes an implementation of Algorithm 1. ZonoOpt is implemented in C++ and Python bindings are available for all classes and functions. Sparse linear algebra is utilized throughout using the Eigen linear algebra library [47]. ZonoOpt is influenced by the MATLAB toolbox zonoLAB [1].
@@ VI-B Model Predictive Control
Consider the MPC problem described in Sec. V-A using the discrete-time double integrator model (35). The time-varying constraint set $\mathcal{S}_{k}$ is decomposed into a time-varying position component and a constant velocity component as $\mathcal{S}_{k}=\mathcal{P}_{k}\times\mathcal{V}$. The velocity and input constraints $\mathcal{V}$ and $\mathcal{U}$ are regular 12-sided polygons, and time-varying position constraints are denoted by $\mathcal{P}_{k}$. In this example, the $\mathcal{P}_{k}$ are given by centrally symmetric hexagons. The regular polygon constraints approximate the quadratic constraints $v_{x,k}^{2}+v_{y,k}^{2}\leq v_{\text{max}}^{2}$ and $a_{x,k}^{2}+a_{y,k}^{2}\leq(v_{\text{min}}\omega_{\text{max}})^{2}$, where $v_{\text{min}}$, $v_{\text{max}}$, and $\omega_{\text{max}}$ are linear and angular velocity limits [48]. For motion planning problems, the time-varying position constraints $\mathcal{P}_{k}$ may be specified by a higher-level autonomous planning algorithm such as in [49].
[…]
Algorithm 2 is used to build the MPC optimization problem, and Algorithm 1 is used to solve it. The optimal trajectory for this example is given in Fig. 4. Note that this trajectory corresponds to a single solution of the MPC optimization problem (i.e., the MPC is not implemented in closed-loop in this example).
@@
Solution times for the MPC optimization problem are compared for the considered QP solvers (ZonoOpt, OSQP, and Gurobi) and problem formulations (Algorithm 2 vs. (36)) in Fig. 5. To assess the scalability of these solution approaches to large MPC horizons, the MPC problem was modified using a scaling factor $f$ such that $N=fN_{0}$ and $\Delta t=\Delta t_{0}/f$ where $N_{0}=55$ and $\Delta t_{0}=1$ s. The position constraint sets $\mathcal{P}_{k}$ and reference states $\mathbf{x}^{r}_{k}$ are correspondingly adjusted via interpolation. The MPC optimization problem was solved for $f=\{1,2,...,21\}$, resulting in MPC horizons ranging from $N=55$ to $N=1155$. Fig. 5 empirically shows that the proposed MPC problem formulation using reachability analysis of constrained zonotopes and corresponding solution approach using ADMM achieves lower solution times and improved scalability with respect to MPC horizon when compared to state-of-the-art problem formulations and QP solvers.

### v3

@@ III-B Numerical Example: Reachable Sets of a Second-Order System
Fig. 2 shows the reachable sets $\mathcal{X}_{k}$ for this system for $k\in\{0,...,N\}$ where $N=15$. The sparsity patterns for the resulting constrained zonotope generator and constraint matrices $G_{N}$ and $A_{N}$ used to represent $\mathcal{X}_{N}$ are shown in Fig. 3. The number of non-zero elements in $G_{N}$ and $A_{N}$ are given explicitly in Table II. Reachability calculations using Eq. (18) result in the constrained zonotope with the fewest non-zero elements in both $G_{N}$ and $A_{N}$. Further, this approach is tied with (11) for producing a constrained zonotope with the fewest number of generators $n_{G}$ and constraints $n_{C}$.
@@
Figure 3: Sparsity plots for matrices $G_{N}$ and $A_{N}$ where $\mathcal{X}_{N}=\left\langle G_{N},\mathbf{c}_{N},A_{N},\mathbf{b}_{N}\right\rangle$ using the standard (11), graph of function (15), and sparsity-promoting (18) reachability calculations to build the $N$-step reachable set $\mathcal{X}_{N}$.
TABLE II: Number of non-zero elements in the generator matrix $G_{N}$ and constraint matrix $A_{N}$ for $\mathcal{X}_{N}$.
@@
$\operatorname*{\mathrm{nnz}}{(G_{N})}$ 33 5 2
$\operatorname*{\mathrm{nnz}}{(A_{N})}$ 315 237 105
@@ IV-A ADMM Formulation
it is clear that (22) is a QP with constraints $\bm{\xi}\in\mathcal{A}\cap\mathcal{B}$ where $\mathcal{A}=\left\{\bm{\xi}\in^{n_{G}}\middle|A\bm{\xi}=\mathbf{b}\right\}$ is an affine set and $\mathcal{B}=\left\{\bm{\xi}\in^{n_{G}}\middle|-\mathbf{1}\leq\bm{\xi}\leq\mathbf{1}\right\}$ is a box.
@@
The ADMM iterations (5) are then
$\bm{\xi}_{k+1}=$ $\operatorname*{\mathrm{argmin}}_{\bm{\xi}}\left(\frac{1}{2}\bm{\xi}^{T}(\tilde{P}+\rho I)\bm{\xi}+(\tilde{\mathbf{q}}+\rho(\mathbf{u}_{k}-\bm{\zeta}_{k}))^{T}\bm{\xi}\right)\;,$
$\text{s.t.}~A\bm{\xi}=\mathbf{b}\;,$ (25a)
[…]
$\text{s.t.}~-\mathbf{1}\leq\bm{\zeta}\leq\mathbf{1}\;,$ (25b)
$\mathbf{u}_{k+1}=$ $\mathbf{u}_{k}+\bm{\xi}_{k+1}-\bm{\zeta}_{k+1}\;.$ (25c)
Eq. (25a) is an equality-constrained quadratic optimization problem and (25b) is a projection onto a box, so the iterations reduce to
@@
where $\min(\cdot,\cdot)$ and $\max(\cdot,\cdot)$ denote element-wise operations, and
$M=\begin{bmatrix}\tilde{P}+\rho I&A^{T}\\ A&0\end{bmatrix}\;$. (27)
@@
If Assumption 1 does not hold, then redundant constraints can be removed as in [54, Alg. 1] or using QR decomposition as in [6].
@@
The ADMM formulation in (26) exploits the structure of the constrained zonotope set representation. As an illustrative example, consider the case where $\mathcal{Z}$ is a centrally symmetric hexagon $\mathcal{Z}=\left\langle G=\begin{bmatrix}\mathbf{g}_{1}&\mathbf{g}_{2}&\mathbf{g}_{3}\end{bmatrix},\mathbf{c},[],[]\right\rangle$ and $P=I$. The matrix $M$ then becomes $M=G^{T}G+\rho I$, which has up to $9$ non-zero elements. In the sparsity-promoting iterations presented in Sec. III, much of the set complexity is moved to the constraint matrix $A$, so we additionally consider the representation $\mathcal{Z}=\left\langle\begin{bmatrix}I&0\end{bmatrix},\mathbf{0},\begin{bmatrix}I&-G\end{bmatrix},\mathbf{0}\right\rangle$. In this case, $M$ will have up to 21 non-zero elements. Representing the same set using H-rep polytopes requires $6$ inequality constraints $\mathbf{a}_{i}^{T}\mathbf{x}\leq b_{i}$. The addition of $6$ slack variables $s_{i}>0$ transforms the inequality constraints into equality constraints such that the problem can be solved via a more general version of the ADMM formulation given in (26) (see [28, 29, 30]). The resulting $A$ matrix has up to $18$ non-zero elements in this case and $P+\rho I$ has $8$ non-zero elements due to the addition of the slack variables, so $M$ will have up to $44$ non-zero elements using H-rep.
Note that $M$ will not have fewer non-zero elements using a constrained zonotope representation $\mathcal{Z}$ vs. H-rep in general. In practice, $M$ has fewer non-zero elements when 1) sparsity-promoting operations are used to compute $\mathcal{Z}$ (Sec. III), and 2) the number of equality constraints in $\mathcal{Z}$ is minimized (e.g., by using operations on zonotopes to build $\mathcal{Z}$ when possible).
The ADMM formulation in (26) is more concise for optimizing over constrained zonotopes than other ADMM formulations. For example, the widely used ADMM solver OSQP [52] solves QPs with constraints of the form $C\mathbf{x}=\mathbf{z}$, $\mathbf{z}_{l}\leq\mathbf{z}\leq\mathbf{z}_{u}$. To optimize over $\mathcal{Z}$ using this form requires $C=\begin{bmatrix}A^{T}&I^{T}\end{bmatrix}^{T}$. This adds $2n_{G}$ non-zero elements to the linear system that OSQP must solve and adds $n_{C}$ projection operations.
Combining constrained zonotopes with high sparsity and the ADMM iterations (26) results in an approach to solving QPs that scales well to large problem sizes.
@@ IV-B Infeasibility Detection
In the following theorem, we propose an infeasibility certificate for (24) that is exact within machine precision and does not require any user-defined tolerances. The interval notation $\bm{\beta}]$ is used to denote the set $\mathcal{B}=\left\{\bm{\xi}\in^{n_{G}}\middle|-\mathbf{1}\leq\bm{\xi}\leq\mathbf{1}\right\}$.
@@
Problem (24) is infeasible if $\exists\mathbf{v}\in\text{span}(A^{T})$ and $\exists\tilde{\bm{\xi}}\in\mathcal{A}$ such that $\mathbf{v}^{T}\tilde{\bm{\xi}}\notin\mathbf{v}^{T}[\bm{\beta}]$.
@@
If (24) is infeasible and $Z^{T}\tilde{P}Z$ is positive definite where $Z$ is a basis for the null space of AA, then the difference of the iterates $\bm{\zeta}_{k}-\bm{\xi}_{k}{=}\mathbf{v}$ for $k\geq k^{\prime}$ with $k^{\prime}$ sufficiently large, where $\mathbf{v}$ is an infeasibility certificate for problem (24).
@@
To implement Thm. 1 for infeasibility detection, the difference of the iterates $\bm{\zeta}_{k}-\bm{\xi}_{k}$ is projected onto the span of $A^{T}$ to generate a candidate infeasibility certificate vector $\mathbf{v}$. If $\mathbf{v}$ satisfies Thm. 1, then (24) is infeasible regardless of whether $k\geq k^{\prime}$ as defined in Cor. 1. In practice, this results in infeasibility certificates being generated in as few as $k=1$ iterations for many problems. Empirically, infeasibility certificates are produced using this implementation for $Z^{T}\tilde{P}Z$ positive semi-definite as well.
@@ IV-C Interpretation of ADMM Iterations
Consider an MPC optimization problem for the system (10) for which the feasible set $\mathcal{Z}$ is defined such that any point $\mathbf{z}\in\mathcal{Z}$ satisfies
$\mathbf{z}=\begin{bmatrix}\mathbf{x}_{0}^{T}&\mathbf{u}_{0}^{T}&\cdots&\mathbf{u}_{N-1}^{T}&\mathbf{x}_{N}^{T}\end{bmatrix}^{T}\;,$ (28a)
$\mathbf{x}_{0}\in\mathcal{X}_{0}\;,$ (28b)
$\forall k\in\{0,..,N-1\}:$
$\hphantom{\forall k}\mathbf{x}_{k}\in\mathcal{S},\;\mathbf{u}_{k}\in\mathcal{U},\;\mathbf{x}_{k+1}\in\{A\mathbf{x}_{k}+B\mathbf{u}_{k}\}\cap\mathcal{S}\;.$ (28c)
@@
In [50], an MPC formulation using ADMM with early termination was presented that decoupled enforcement of system dynamics constraints from state and input feasibility constraints between ADMM steps. In that formulation, the equality constraints enforced in (5a) correspond to the system dynamics while feasibility constraints are enforced in (5b). This is analogous to building the MPC problem using (18) with zonotopes in our framework. Using (11) with zonotopes to build the MPC problem would instead simultaneously enforce dynamics and input constraints with state constraints decoupled.
@@ IV-D ADMM Algorithm for Constrained Zonotopes
Inputs: $\mathcal{Z}$, $P$, $\mathbf{q}$, $\rho$, $\epsilon_{p}$, $\epsilon_{d}$
Outputs: $\mathbf{x}^{*}\in\mathcal{Z}$, infeasibility certificate flag
@@ V Applications
The following application examples demonstrate the utility of combining the sparse reachability operations presented in Sec. III with the ADMM algorithm presented in Sec. IV to solve control, estimation, and safety verification problems. In all of the presented examples, efficient problem formulations rely on the sparsity-promoting iterations (18).
@@ V-A Model Predictive Control
${\mathbf{x}_{k}\in\mathcal{S}_{k}\;,\;\mathbf{x}_{N}\in\mathcal{S}_{N}}\;,\;\mathbf{u}_{k}\in\mathcal{U}\;,$ (29c)
@@
Inputs: $\mathbf{x}_{0}$, $\mathbf{x}^{r}_{k}$, $Q$, $R$, $Q_{N}$, $A$, $B$, NN
Outputs: $\mathcal{Z},P,\mathbf{q}$
@@
13:return $(\mathcal{Z},P,\mathbf{q})$
@@ V-B Moving Horizon Estimation and Set-Valued State Estimation
Assuming $\mathcal{S}$, $\mathcal{W}$, $\mathcal{V}$, and $\mathcal{X}_{-N}$ are zonotopes or constrained zonotopes, reachability analysis with constrained zonotopes can be used to solve the SVSE problem for (30) via the iterations
@@
The process and measurement noise covariance matrices are $Q$ and $R$, respectively. The covariance of the state estimate at time step k=Nk=-N is $P_{-N}$. The prior state estimate for k=Nk=-N is $\hat{\mathbf{x}}_{-N}$. The set of possible states at time step k=Nk=-N is $\mathcal{X}_{-N}$, which is updated recursively between MHE iterations by applying (32).
Algorithm 3 shows how the MHE optimization problem (33) can be written in the form of (21) using constrained zonotopes and the set operations in (3). Line 5 implements (32) without projecting onto $\mathcal{X}_{k+1}$. Elements $\mathbf{z}\in\mathcal{Z}$ where $\mathcal{Z}$ is the set of feasible solutions to (32) have the structure
$\mathbf{z}=\begin{bmatrix}\mathbf{x}_{-N}^{T}&\mathbf{w}_{-N}^{T}&\cdots&\mathbf{w}_{-1}^{T}&\mathbf{x}_{0}^{T}\end{bmatrix}\;$. (34)
As such, the solution $\mathcal{X}_{0}$ to the SVSE problem at time step $k=0$ can be readily extracted via a linear map operation. After solving (21), a best state estimate $\hat{\mathbf{x}}_{0}\in\mathcal{X}_{0}$ can also be extracted.
@@
Inputs: $\mathcal{X}_{-N}$, $\mathcal{W}$, $\mathcal{V}$, $\hat{\mathbf{x}}_{-N}$, $\mathbf{y}_{k}$, $\mathbf{u}_{k}$, $P^{-1}_{-N}$, $A$, $B$, $C$, $Q$, $R$, NN
Outputs: $\mathcal{Z}$, $P$, $\mathbf{q}$, $\mathcal{X}_{0}$
@@ VI-A Implementation
All examples are implemented using the open-source toolbox ZonoOpt , which was developed as a contribution of this paper. This toolbox provides classes for zonotopes, constrained zonotopes, and hybrid zonotopes, implements the set operations in (3) (among others), and includes an implementation of Algorithm 1. ZonoOpt is implemented in C++ and Python bindings are available for all classes and functions. Sparse linear algebra is utilized throughout using the Eigen linear algebra library [15]. ZonoOpt is influenced by the MATLAB toolbox zonoLAB [21].
@@ VI-B Model Predictive Control
Consider the MPC problem described in Sec. V-A using the discrete-time double integrator model (37). This problem is depicted in Fig. 4. The time-varying constraint set $\mathcal{S}_{k}$ is decomposed into a time-varying position component and a constant velocity component as $\mathcal{S}_{k}=\mathcal{P}_{k}\times\mathcal{V}$. The velocity and input constraints $\mathcal{V}$ and $\mathcal{U}$ are regular 12-sided polygons, and time-varying position constraints are denoted by $\mathcal{P}_{k}$. In this example, the $\mathcal{P}_{k}$ are given by centrally symmetric hexagons. The $\mathcal{P}_{k}$ define a corridor that the system must remain inside, and the position references $\mathbf{x}_{k}^{r}$ are centered within this corridor. For motion planning problems, the $\mathcal{P}_{k}$ may be specified by a higher-level autonomous planning algorithm such as in [57]. The regular polygon constraints approximate the quadratic constraints $v_{x,k}^{2}+v_{y,k}^{2}\leq v_{\text{max}}^{2}$ and $a_{x,k}^{2}+a_{y,k}^{2}\leq(v_{\text{min}}\omega_{\text{max}})^{2}$, where $v_{\text{min}}$, $v_{\text{max}}$, and $\omega_{\text{max}}$ are linear and angular velocity limits [56].
[…]
Algorithm 2 is used to build the MPC optimization problem, and Algorithm 1 is used to solve it. The optimal trajectory for this example is depicted as red dots in Fig. 4. Note that this trajectory corresponds to a single solution of the MPC optimization problem (i.e., the MPC is not implemented in closed loop in this example).
@@
ZonoOpt
CZ
OSQP
CZ
OSQP
H-rep
Gurobi
CZ
Gurobi
H-rep
$10^{0}$$10^{1}$$10^{2}$Solution time [ms]
Figure 6: MPC solution time statistics for closed-loop simulation. In this box and whisker plot, the medians are depicted with black lines, the boxes enclose the middle 50% of data points, and the whiskers enclose all data points.
@@
When Algorithm 2 is used to build the MPC problem, the $M$ matrix (27) has $12315$ non-zero elements and there are $825$ optimization variables $\bm{\xi}$. Using (38) to build the MPC problem, an equivalent ADMM formulation would require an $M$ matrix with $13052$ elements and $1984$ optimization variables. See Sec. IV-A for further discussion of ADMM complexity using a constrained zonotope constraint representation.
Solution times for the MPC optimization problem are compared for the considered QP solvers (ZonoOpt, OSQP, and Gurobi) and problem formulations (Algorithm 2 vs. (38)) in Fig. 5. To assess the scalability of these solution approaches to large MPC horizons, the MPC problem was modified using a scaling factor $f$ such that $N=fN_{0}$ and $\Delta t=\Delta t_{0}/f$ where $N_{0}=55$ and $\Delta t_{0}=1$ s. The position constraint sets $\mathcal{P}_{k}$ and reference states $\mathbf{x}^{r}_{k}$ are correspondingly adjusted via interpolation. The MPC optimization problem was solved for $f=\{1,2,...,21\}$, resulting in MPC horizons ranging from $N=55$ to $N=1155$.
We additionally simulate the MPC in closed loop for the problem depicted in Fig. 4. Here, the scaling factor was set to $f=4$ and the system was simulated for $220$ time steps. The receding horizon length was $N=100$. MPC solution time statistics over the course of the simulation for each solution approach are shown in Fig. 6.
Figs. 5 and 6 empirically show that the proposed MPC problem formulation using reachability analysis of constrained zonotopes and corresponding solution approach using ADMM achieves lower solution times on average and improved scalability with respect to MPC horizon when compared to state-of-the-art problem formulations and QP solvers. These efficiency improvements are a result of the efficient use of the constrained zonotope structure in the ADMM iterations.